Circular No. 8572

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2002hh AND 2005V R. J. Beswick, D. Fenech, H. Thrall, M. K. Argo, T. W. B. Muxlow,

and A. Pedlar, Jodrell Bank Observatory, report further detections of the radio emission from the evolving type-II supernova 2002hh in NGC 6946 (IAUC 8005, 8018) and limits on the radio emission from SN 2005V in NGC 2146. Very Large Array ‘B’-configuration observations of 2002hh detected flux densities of 0.35 ± 0.1 mJy at 4.860 GHz on 2003 Nov. 14 and 1.6 ± 0.2 mJy at 1.425 GHz on 2005 Apr. 16 at α = 20^{h}34^{m}44^{s}.25, δ = +60^{o}07′19′′.4 (equinox 2000.0; ± 0′′.2 in each coordinate), in close agreement with previously reported positions. MERLIN observations place 3σ upper limits for the type-Ib/c supernova 2005V (cf. IAUC 8474) on 2005 Apr. 4 of 0.14 mJy at 1.658 GHz, and on May 19 of 0.23 mJy at 1.408 GHz at the reported position. Further observations of both sources are underway.

COMETS C/2005 L14, C/2005 L15, C/2005 M2–M5 (SOHO) Five more tailless Kreutz sungrazing comets, and one elongated ‘non-

group’ comet (C/2005 M3), have been found on SOHO images (cf. IAUC 8570; JZ = J. Zhang). C/2005 L14 was tiny and at the background noise level. C/2005 L15 was stellar and relatively bright, reaching mag ∼ 5.0 at 9.1R_{⊙} on June 16.512 UT. C/2005 M2 appeared very small and stellar in C3 images but slightly diffuse in C2 images, reaching mag ∼ 6.8 at 7.6R_{⊙} on June 16.592. C/2005 M3 reached mag ∼ 7.6 at 7.3R_{⊙} on June 19.700. C/2005 M4 was diffuse, reaching mag ∼ 7.7 at 7.9R_{⊙} on June 20.783. C/2005 M5 was very diffuse, reaching mag ∼ 8.2 at 7.6R_{⊙} on June 21.033.

Comet 2005 UT α_{2000} δ_{2000} Inst. F MPEC
C/2005 L14 June 12.558 5^{h}19.^{m}8 +21^{o}18′ C2  TH 2005-O26
C/2005 L15 15.779 5 34.3 +19 24 C3 XL 2005-O26
C/2005 M2  16.238 5 34.5 +20 35 C3/2 KB 2005-O26
C/2005 M3  19.617 5 46.0 +22 10 C2  HS 2005-O26
C/2005 M4  20.750 5 51.9 +21 35 C2  JZ 2005-O26
C/2005 M5  20.975 5 53.3 +21 33 C2  TH 2005-O26

COMET 9P/TEMPEL Further to IAUC 8558, J. McGaha revises his report concerning images

obtained with the 0.62-m f/5.1 reflector to note the following increases in R magnitude (estimated uncertainty ± 0.05 mag) seen in the stated photometric aperture sizes: 3′′, 1.48; 10′′, 1.33; 16′′, 0.99; 32′′, 0.64.

2005 July 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green